# Asymptotic behaviour of the Hitchin metric on the moduli space of Higgs bundles

### v2

@@ 2.5.2 Universal boundedness of higher derivatives of Higgs fields
Let $0<r_{1}<r_{0}$. We set $r_{2}=\frac{r_{0}+r_{1}}{2}$. By a variant of Simpson’s main estimate (see [20, Proposition 2.1]), we obtain the following estimate in [23, Proposition 4.1].
@@ 2.5.3 The sup norm and the $L^{2}$-norm of harmonic $1$-forms
Proof   Let $\epsilon_{0}$ be as in Theorem 2.32. Let $B_{2}$ be as in Proposition 2.34. We set $\epsilon_{1}:=\min\{2^{-1}(r_{2}-r_{1}),B_{2}^{-1/2}\epsilon_{0}^{1/2}\}$, which satisfies $0<\epsilon_{1}<r_{2}-r_{1}$ and $B_{2}\epsilon_{1}^{2}\leq\epsilon_{0}$. For any $z_{0}\in\mathcal{B}(r_{1})$, let $\psi_{z_{0}}:\mathcal{B}(1)\longrightarrow\mathcal{B}(r_{2})$ be defined by $\psi_{z_{0}}(\zeta)=z_{0}+\epsilon_{1}\zeta$. We obtain $(\widehat{E}_{z_{0}},\widehat{\theta}_{z_{0}},\widehat{h}_{z_{0}})=\psi_{z_{0}}^{\ast}(E,\theta,h)$ and $\widehat{\tau}_{z_{0}}=\psi_{z_{0}}^{\ast}(\tau)$ on $\mathcal{B}(1)$. Let $\nabla_{\widehat{h}_{z_{0}}}$ denote the Chern connection of $(\widehat{E}_{z_{0}},\widehat{h}_{z_{0}})$, and let $R(\widehat{h}_{z_{0}})$ denote the curvature of $\nabla_{\widehat{h}_{z_{0}}}$. Let $\|\widehat{\tau}_{z_{0}}\|_{L^{2},\widehat{h}_{z_{0}}}$ denote the $L^{2}$-norm of $\widehat{\tau}_{z_{0}}$ with respect to $\widehat{h}_{z_{0}}$ and $g_{0}$. We obtain
@@
By using Theorem 2.32, Lemma 2.33, and the bootstrapping argument, we can show that there exists $B_{5}>0$, depending only on $B_{2}$, such that $|\widehat{\tau}_{z_{0}}|_{\widehat{h}_{z_{0}},g_{0}}\leq B_{5}\|\tau\|_{L^{2},h}$ on $\{|\zeta|<1/4\}$. It implies that $|\tau|_{h,g_{0}}\leq\epsilon_{1}^{-1}B_{5}\|\tau\|_{L^{2},h}$ on $\mathcal{B}(r_{1})$.
@@ 2.5.4 Poisson equation
Proof   We use the notation in the proof of Proposition 2.35. We obtain the section $\widehat{s}_{z_{0}}=\psi_{z_{0}}^{-1}(s)$. By using Theorem 2.32, Lemma 2.33, Proposition 2.34 and the bootstrapping argument, we obtain the estimates for $|\widehat{s}_{z_{0}}|_{\widehat{h}_{z_{0}}}$ and |(¯$|(\overline{\partial}_{\widehat{E}_{z_{0}}}+\widehat{\theta}_{z_{0}})\widehat{s}_{z_{0}}|_{\widehat{h}_{z_{0}},g_{0}}$. Then, we obtain (25).
@@ 2.6.2 Real bilinear forms on complex vector spaces
Let $V$ be a finite dimensional ${\mathbb{C}}$-vector space. Let $V_{{\mathbb{R}}}$ denote the underlying ${\mathbb{R}}$-vector space. We set $V^{\lor}=\mathop{\rm Hom}\nolimits_{{\mathbb{C}}}(V,{\mathbb{C}})$ and $V_{{\mathbb{R}}}^{\lor}=\mathop{\rm Hom}\nolimits_{{\mathbb{R}}}(V_{{\mathbb{R}}},{\mathbb{R}})$. The multiplication of $\sqrt{-1}$ on $V$ induces the ${\mathbb{R}}$-automorphism $I$ on $V_{{\mathbb{R}}}$ such that $I^{2}=-\mathop{\rm id}\nolimits_{V_{{\mathbb{R}}}}$. It induces an ${\mathbb{R}}$-linear automorphism $I^{\lor}$ on $V_{{\mathbb{R}}}^{\lor}$ by $I^{\lor}(f)(v)=f(Iv)$.
@@
where $V^{\lor}$ is the $\sqrt{-1}$-eigen space of $I^{\lor}$, and $\overline{V^{\lor}}$ is the $-\sqrt{-1}$-eigen space of $I^{\lor}$. We obtain the isomorphism of ${\mathbb{R}}$-vector spaces $\Phi:V_{{\mathbb{R}}}^{\lor}\to V^{\lor}$ as the composition of the inclusion $V_{{\mathbb{R}}}^{\lor}\to V_{{\mathbb{R}}}^{\lor}\otimes{\mathbb{C}}$ and the projection $V_{{\mathbb{R}}}^{\lor}\otimes{\mathbb{C}}=V^{\lor}\oplus\overline{V^{\lor}}\to V^{\lor}$. It is explicitly described as
@@
  • Suppose that $I$ is an isometry with respect to $g_{{\mathbb{R}}}$, and that $\omega(u,v)=g_{{\mathbb{R}}}(u,Iv)$ for any $u,v$. Then, we obtain $\omega^{\lor}(u^{\lor},v^{\lor})=-g_{{\mathbb{R}}}^{\lor}(u^{\lor},I^{\lor}v^{\lor})$. In particular, if $\omega(\cdot,I\cdot)$ is positive definite, $-\omega^{\lor}(\cdot,I^{\lor},\cdot)$ is positive definite.
@@ 2.7.1 Bilinear forms on $H^{1}(\Sigma,\mathcal{O}_{\Sigma})$
There exist harmonic representative $\eta_{i}$ of $a_{i}$, i.e., $\eta_{i}$ are $(0,1)$-forms such that $\partial_{\Sigma}\eta_{i}=0$, and $[\eta_{i}]=a_{i}$ in $H^{1}(\Sigma,\mathcal{O}_{\Sigma})$. We have
@@ 2.7.3 Duality
$-\omega_{1}^{\lor}(a_{1},a_{2})=-\omega_{1}\bigl{(}\widetilde{\Psi}_{\omega_{1}}^{-1}(a_{1}),\widetilde{\Psi}_{\omega_{1}}^{-1}(a_{2})\bigr{)}=-\omega_{1}(\overline{a}_{1},\overline{a}_{2})=\int_{\Sigma}\bigl{(}\overline{a}_{1}a_{2}+a_{1}\overline{a}_{2}\bigr{)}=\omega_{2}(a_{1},a_{2}).$
@@ 2.8 Estimate for the family of the associated Laplacian operators
Let $X$ be a compact Riemann surface with a conformal metric $g_{X}$. Let $(E,\theta)$ be a generically regular semisimple Higgs bundle of degree 00 whose spectral curve is smooth. We fix a flat metric $h_{\det(E)}$ of $\det(E)$. For each $t>0$, there exists a harmonic metric $h_{t}$ of $(E,t\theta)$. We recall that the sequence $h_{t}$ is convergent to the limiting configuration $h_{\infty}$ in the $C^{\infty}$-sense locally on $X^{\circ}$.
Let $\mathop{\rm End}\nolimits_{0}(E)$ denote the trace free part of $\mathop{\rm End}\nolimits(E)$. Let (¯$(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{t}}$ denote the formal adjoint of ¯$\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta$ with respect to $h_{t}$ and $g_{X}$. We consider the operators $\Delta_{t}=(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{t}}\circ(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)$ on $A^{0}(\mathop{\rm End}\nolimits_{0}(E))$. For any $s$ of $A^{0}(\mathop{\rm End}\nolimits_{0}(E))$, let $\|s\|_{L^{2},h_{t}}$ denote the $L^{2}$-norm of $s$ with respect to $h_{t}$ and $g_{X}$.
@@
Proof   If $s\in A^{0}(\mathop{\rm End}\nolimits_{0}(E))$ satisfies $\Delta_{t}s=0$, it is standard to obtain (¯$(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)s=0$. Because $(E,t\theta)$ is stable, we obtain $s=0$.
@@
By taking a subsequence, we may assume that the sequence $s_{i}$ is weakly convergent in $L_{1}^{2}$ locally on $X^{\circ}$. Let $s_{\infty}$ denote the limit. The sequence $s_{i}$ $(i=1,2,\ldots)$ is convergent to $s_{\infty}$ in $L^{p}$ for any $p>1$ locally on $X^{\circ}$. We obtain $\theta(s_{\infty})=0$ and ¯$\overline{\partial}_{E}(s_{\infty})=0$. We also have $\|s_{\infty}\|_{L^{2},h_{\infty}}\leq 1$. Because $\mathop{\rm tr}\nolimits s_{i}=0$, we obtain $\mathop{\rm tr}\nolimits s_{\infty}=0$.
@@
Let $U$ be a relatively compact neighbourhood of $P_{\infty}$ in $X$ with a holomorphic coordinate $z$. Let (¯$(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{i}}$ denote the formal adjoint of ¯$\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta$ with respect to $h_{i}$ and the Euclidean metric $dz\,d\overline{z}$. By Lemma 2.45 below, we obtain
$-\partial_{z}\partial_{\overline{z}}|\widetilde{s}_{i}|_{h_{i}}^{2}\leq 2\bigl{|}(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{i}}(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)(\widetilde{s}_{i})\bigr{|}_{h_{i}}\cdot|\widetilde{s}_{i}|_{h_{i}}=2a_{t(i)}g_{U}|\widetilde{s}_{i}|_{h_{i}}^{2}\leq 2a_{t(i)}g_{U}.$
@@
$-\partial_{z}\partial_{\overline{z}}\bigl{(}|\widetilde{s}_{i}|_{h_{i}}^{2}-2a_{t(i)}\beta_{U}\bigr{)}\leq 0.$
@@
Let us consider the case $P_{\infty}\in D(\theta)$. Let $U^{\prime}$ be a relatively compact neighbourhood of $P_{\infty}$ in $U$ such that $U^{\prime}\cap D(\theta)=\{P_{\infty}\}$. By applying the maximum principle to the subharmonic function $|\widetilde{s}_{i}|_{h_{i}}^{2}-2a_{t(i)}\beta_{U}$, there exists $Q_{i}\in\partial U^{\prime}$ such that
@@
$-\partial_{z}\partial_{\overline{z}}|s|_{h}^{2}=2\mathop{\rm Re}\nolimits\bigl{(}(-\nabla_{z}\nabla_{\overline{z}}+f_{h}^{\dagger}\circ f)s,s\bigr{)}-|\nabla_{z}s|_{h}^{2}-|\nabla_{\overline{z}}s|_{h}^{2}-|f(s)|_{h}^{2}-|f^{\dagger}_{h}(s)|_{h}^{2}.$ (28)
@@
$-\partial_{z}\partial_{\overline{z}}|s|_{h}^{2}=-h\bigl{(}\nabla_{z}\nabla_{\overline{z}}s,s\bigr{)}-h\bigl{(}s,\nabla_{\overline{z}}\nabla_{z}s\bigr{)}-\bigl{|}\nabla_{\overline{z}}s\bigr{|}_{h}^{2}-\bigl{|}\nabla_{z}s\bigr{|}_{h}^{2}\\ =-h\bigl{(}\nabla_{z}\nabla_{\overline{z}}s,s\bigr{)}-h\bigl{(}s,\nabla_{z}\nabla_{\overline{z}}s\bigr{)}-h\bigl{(}s,[f,f^{\dagger}_{h}]s\bigr{)}-\bigl{|}\nabla_{\overline{z}}s\bigr{|}_{h}^{2}-\bigl{|}\nabla_{z}s\bigr{|}_{h}^{2}\\ =2\mathop{\rm Re}\nolimits h\bigl{(}-\nabla_{z}\nabla_{\overline{z}}s,s\bigr{)}+2h\bigl{(}s,f^{\dagger}_{h}\circ f(s)\bigr{)}-\bigl{|}f(s)\bigr{|}_{h}^{2}-\bigl{|}f^{\dagger}_{h}(s)\bigr{|}_{h}^{2}-\bigl{|}\nabla_{\overline{z}}s\bigr{|}_{h}^{2}-\bigl{|}\nabla_{z}s\bigr{|}_{h}^{2}\\ =2\mathop{\rm Re}\nolimits h\bigl{(}(-\nabla_{z}\nabla_{\overline{z}}+f^{\dagger}_{h}f)s,s\bigr{)}-\bigl{|}f(s)\bigr{|}_{h}^{2}-\bigl{|}f^{\dagger}_{h}(s)\bigr{|}_{h}^{2}-\bigl{|}\nabla_{\overline{z}}s\bigr{|}_{h}^{2}-\bigl{|}\nabla_{z}s\bigr{|}_{h}^{2}.$ (29)

### v3

@@ 2.5.2 Universal boundedness of higher derivatives of Higgs fields
Let $0<r_{1}<r_{0}$. We set $r_{2}=\frac{r_{0}+r_{1}}{2}$. By a variant of Simpson’s main estimate (see [20, Proposition 2.1]), we obtain the following estimate in [24, Proposition 4.1].
@@ 2.5.3 The sup norm and the $L^{2}$-norm of harmonic $1$-forms
Proof  Let $\epsilon_{0}$ be as in Theorem 2.32. Let $B_{2}$ be as in Proposition 2.33. We set $\epsilon_{1}:=\min\{2^{-1}(r_{2}-r_{1}),B_{2}^{-1/2}\epsilon_{0}^{1/2}\}$, which satisfies $0<\epsilon_{1}<r_{2}-r_{1}$ and $B_{2}\epsilon_{1}^{2}\leq\epsilon_{0}$. For any $z_{0}\in\mathcal{B}(r_{1})$, let $\psi_{z_{0}}:\mathcal{B}(1)\longrightarrow\mathcal{B}(r_{2})$ be defined by $\psi_{z_{0}}(\zeta)=z_{0}+\epsilon_{1}\zeta$. We obtain $(\widehat{E}_{z_{0}},\widehat{\theta}_{z_{0}},\widehat{h}_{z_{0}})=\psi_{z_{0}}^{\ast}(E,\theta,h)$ and $\widehat{\tau}_{z_{0}}=\psi_{z_{0}}^{\ast}(\tau)$ on $\mathcal{B}(1)$. Let $\nabla_{\widehat{h}_{z_{0}}}$ denote the Chern connection of $(\widehat{E}_{z_{0}},\widehat{h}_{z_{0}})$, and let $R(\widehat{h}_{z_{0}})$ denote the curvature of $\nabla_{\widehat{h}_{z_{0}}}$. Let $\|\widehat{\tau}_{z_{0}}\|_{L^{2},\widehat{h}_{z_{0}}}$ denote the $L^{2}$-norm of $\widehat{\tau}_{z_{0}}$ with respect to $\widehat{h}_{z_{0}}$ and $g_{0}$. We obtain the following on $\mathcal{B}(1)$:
@@
Let $\boldsymbol{v}$} be an orthonormal frame of $(\widehat{E}_{z_{0}},\widehat{h}_{z_{0}})$ as in Theorem 2.32. Let $\mathcal{A}$ be the connection form of $\nabla_{\widehat{h}_{z_{0}}}$ with respect to $\boldsymbol{v}$}. The $(p,q)$-part is denoted by $\mathcal{A}^{p,q}$. Let $\Theta$ denote the matrix valued $(1,0)$-form determined by $\widehat{\theta}_{z_{0}}{\boldsymbol{v}}={\boldsymbol{v}}\Theta$. We have the expression $\tau=\sum\tau_{i}v_{i}$, and we obtain the vector-valued $1$-form $\boldsymbol{\tau}_{z_{0}}=(\tau_{i})$. The $(p,q)$-part of $\boldsymbol{\tau}_{z_{0}}$ is denoted by $\boldsymbol{\tau}_{z_{0}}^{p,q}$. Because $\tau$ is harmonic, we have
$\overline{\partial}\boldsymbol{\tau}_{z_{0}}^{1,0}+\mathcal{A}^{0,1}(\boldsymbol{\tau}_{z_{0}}^{1,0})+\Theta(\boldsymbol{\tau}_{z_{0}}^{0,1})=0,\quad\partial\boldsymbol{\tau}_{z_{0}}^{0,1}+\mathcal{A}^{1,0}(\boldsymbol{\tau}_{z_{0}}^{0,1})+{}^{t}\!\overline{\Theta}(\boldsymbol{\tau}_{z_{0}}^{1,0})=0.$ (24)
By (24), the $L_{1}^{2}$-norm of $\boldsymbol{\tau}_{z_{0}}$ on $\mathcal{B}(2/3)$ is dominated by $\|\tau\|_{L^{2},h}$. It implies that for $p>2$, the $L^{p}$-norm of $\boldsymbol{\tau}_{z_{0}}$ on $\mathcal{B}(1/2)$ is dominated by $\|\tau\|_{L^{2},h}$. By (24) again, the $L_{1}^{p}$-norm of $\boldsymbol{\tau}_{z_{0}}$ on $\mathcal{B}(1/3)$ is dominated by $\|\tau\|_{L^{2},h}$. Hence, we can show that there exists $B_{5}>0$, depending only on $B_{2}$, such that $|\widehat{\tau}_{z_{0}}|_{\widehat{h}_{z_{0}},g_{0}}\leq B_{5}\|\tau\|_{L^{2},h}$ on $\{|\zeta|<1/4\}$. It implies that $|\tau|_{h,g_{0}}\leq\epsilon_{1}^{-1}B_{5}\|\tau\|_{L^{2},h}$ on $\mathcal{B}(r_{1})$.
@@ 2.5.4 Poisson equation
Proof  We set $\sigma=(\overline{\partial}_{E}+\theta)s$ and $\rho=(\partial_{E,h}+\theta^{\dagger})(\overline{\partial}_{E}+\theta)s$. We use the notation in the proof of Proposition 2.34. We obtain the section $s_{z_{0}}=\psi_{z_{0}}^{-1}(s)$, the $1$-form $\sigma_{z_{0}}=\psi_{z_{0}}^{-1}(\sigma)$ and the $2$-form $\rho_{z_{0}}=\psi_{z_{0}}^{-1}(\sigma)$ on $\mathcal{B}(1)$. Let $\kappa:{\mathbb{R}}\to\{0\leq a\leq 1\}$ be a $C^{\infty}$-function such that (i) $\kappa(t)=1$ $(t\leq 3/4)$ and $\kappa(t)=0$ $(t\geq 4/5)$, (ii) $\kappa^{\prime}/\kappa^{1/2}$ is bounded on $\{t\,|\,\kappa(t)>0\}$. Let $\chi(\zeta)=\kappa(|\zeta|)$. We obtain
$\int_{\mathcal{B}(1)}\chi|\sigma_{z_{0}}|^{2}_{\widehat{h}_{z_{0}},g_{0}}\leq\int_{\mathcal{B}(1)}|\overline{\partial}\chi|_{g_{0}}\cdot|s_{z_{0}}|_{\widehat{h}_{z_{0}}}\cdot|\sigma_{z_{0}}|_{\widehat{h}_{z_{0}},g_{0}}+\int_{\mathcal{B}(1)}\chi|s_{z_{0}}|_{\widehat{h}_{z_{0}}}\cdot|\rho_{z_{0}}|_{\widehat{h}_{z_{0}},g_{0}}\\ \leq\left(\int_{\mathcal{B}(1)}|\chi^{-1/2}\overline{\partial}\chi|^{2}|s_{z_{0}}|^{2}_{\widehat{h}_{z_{0}}}\right)^{1/2}\left(\int_{\mathcal{B}(1)}\chi|\sigma_{z_{0}}|^{2}_{\widehat{h}_{z_{0}},g_{0}}\right)^{1/2}+\left(\int_{\mathcal{B}(1)}|s_{z_{0}}|_{\widehat{h}_{z_{0}}}^{2}\right)^{1/2}\cdot\left(\int_{\mathcal{B}(1)}|\rho_{z_{0}}|_{\widehat{h}_{z_{0}},g_{0}}^{2}\right)^{1/2}.$ (26)
There exist $C_{1},C_{2}>0$, depending only on $\kappa$, such that
$\left(\int_{\mathcal{B}(3/4)}|\sigma_{z_{0}}|^{2}_{\widehat{h}_{z_{0}},g_{0}}\right)^{1/2}\leq\left(\int_{\mathcal{B}(1)}\chi|\sigma_{z_{0}}|^{2}_{\widehat{h}_{z_{0}},g_{0}}\right)^{1/2}\leq C_{1}\left(\int_{\mathcal{B}(1)}|s_{z_{0}}|^{2}_{\widehat{h}_{z_{0}}}\right)^{1/2}+C_{2}\left(\int_{\mathcal{B}(1)}|\rho_{z_{0}}|^{2}_{\widehat{h}_{z_{0}},g_{0}}\right)^{1/2}.$ (27)
We have the expressions $s_{z_{0}}=\sum s_{z_{0},i}v_{i}$, $\sigma_{z_{0}}=\sum\sigma_{z_{0},i}v_{i}$ and $\rho_{z_{0}}=\sum\rho_{z_{0},i}v_{i}$. We set ${\boldsymbol{s}}=(s_{i})$, ${\boldsymbol{\sigma}}_{z_{0}}=(\sigma_{z_{0},i})$, and ${\boldsymbol{\rho}}_{z_{0}}=(\rho_{z_{0},i})$. We have
$(\overline{\partial}+\mathcal{A}^{0,1}){\boldsymbol{s}}_{z_{0}}+\Theta{\boldsymbol{s}}={\boldsymbol{\sigma}}_{z_{0}},$ (28)
$(\partial+\mathcal{A}^{1,0}){\boldsymbol{\sigma}}_{z_{0}}^{0,1}+{}^{t}\!\overline{\Theta}({\boldsymbol{\sigma}}^{1,0}_{z_{0}})={\boldsymbol{\rho}}_{z_{0}},\quad(\overline{\partial}+\mathcal{A}^{0,1}){\boldsymbol{\sigma}}_{z_{0}}^{1,0}+\Theta({\boldsymbol{\sigma}}_{z_{0}}^{0,1})=0.$ (29)
By (27), (29) and the elliptic regularity, the $L_{1}^{2}$-norm of ${\boldsymbol{\rho}}_{z_{0}}$ on $\mathcal{B}(2/3)$ is dominated by $\|s\|_{L^{2},h,g_{0}}$ and $\sup|\rho_{z_{0}}|_{\widehat{h}_{0},g_{0}}$. It implies that for $p>2$, the $L^{p}$-norm of ${\boldsymbol{\rho}}_{z_{0}}$ on $\mathcal{B}(1/2)$ is dominated by $\|s\|_{L^{2},h,g_{0}}$ and $\sup|\rho_{z_{0}}|_{\widehat{h}_{0},g_{0}}$. By (27), (29) and the elliptic regularity again, the $L_{1}^{p}$-norm of ${\boldsymbol{\rho}}_{z_{0}}$ on $\mathcal{B}(1/3)$ is dominated by $\|s\|_{L^{2},h,g_{0}}$ and $\sup|\rho_{z_{0}}|_{\widehat{h}_{0},g_{0}}$. Hence, $|\sigma_{z_{0}}|_{\widehat{h}_{0},g_{0}}$ on $\mathcal{B}(1/4)$ is dominated by $\|s\|_{L^{2},h,g_{0}}$ and $\sup|\rho_{z_{0}}|_{\widehat{h}_{0},g_{0}}$. By (28) and the elliptic regularity, $|s_{z_{0}}|$ on $\mathcal{B}(1/5)$ is dominated by $\|s\|_{L^{2},h,g_{0}}$ and $\sup|\rho_{z_{0}}|_{\widehat{h}_{0},g_{0}}$. Then, we obtain (25).
@@ 2.6.2 Real bilinear forms on complex vector spaces
Let $V$ be a finite dimensional $\mathbb{C}$}-vector space. Let $V_{{\mathbb{R}}}$ denote the underlying $\mathbb{R}$}-vector space. We set $V^{\lor}=\mathop{\rm Hom}\nolimits_{{\mathbb{C}}}(V,{\mathbb{C}})$ and $V_{{\mathbb{R}}}^{\lor}=\mathop{\rm Hom}\nolimits_{{\mathbb{R}}}(V_{{\mathbb{R}}},{\mathbb{R}})$. The multiplication of $\sqrt{-1}$ on $V$ induces the $\mathbb{R}$}-automorphism II on $V_{{\mathbb{R}}}$ such that $I^{2}=-\mathop{\rm id}\nolimits_{V_{{\mathbb{R}}}}$. It induces an $\mathbb{R}$}-linear automorphism $I^{\lor}$ on $V_{{\mathbb{R}}}^{\lor}$ by $I^{\lor}(f)(v)=f(Iv)$.
@@
where $V^{\lor}$ is the $\sqrt{-1}$-eigen space of $I^{\lor}$, and $\overline{V^{\lor}}$ is the $\sqrt{-1}-$eigen space of $I^{\lor}$. We obtain the isomorphism of $\mathbb{R}$}-vector spaces $\Phi:V_{{\mathbb{R}}}^{\lor}\to V^{\lor}$ as the composition of the inclusion $V_{{\mathbb{R}}}^{\lor}\to V_{{\mathbb{R}}}^{\lor}\otimes{\mathbb{C}}$ and the projection $V_{{\mathbb{R}}}^{\lor}\otimes{\mathbb{C}}=V^{\lor}\oplus\overline{V^{\lor}}\to V^{\lor}$. It is explicitly described as
@@
  • Suppose that $I$ is an isometry with respect to $g_{{\mathbb{R}}}$, and that $\omega(u,v)=g_{{\mathbb{R}}}(u,Iv)$ for any $u,v$. Then, we obtain $\omega^{\lor}(u^{\lor},v^{\lor})=-g_{{\mathbb{R}}}^{\lor}(u^{\lor},I^{\lor}v^{\lor})$. In particular, if $\omega(I\cdot,\cdot)$ is positive definite, $\omega^{\lor}(I^{\lor}\cdot,\cdot)$ is positive definite.
@@ 2.7.1 Bilinear forms on $H^{1}(\Sigma,\mathcal{O}_{\Sigma})$
There exist the harmonic representatives $\eta_{i}$ of $a_{i}$, i.e., $\eta_{i}$ are the $(0,1)$-forms such that $\partial_{\Sigma}\eta_{i}=0$, and $\eta_{i}]=a_{i}$ in $H^{1}(\Sigma,\mathcal{O}_{\Sigma})$. We have
@@ 2.7.3 Duality
$\omega_{1}^{\lor}(a_{1},a_{2})=-\omega_{1}\bigl(\widetilde{\Psi}_{\omega_{1}}^{-1}(a_{1}),\widetilde{\Psi}_{\omega_{1}}^{-1}(a_{2})\bigr)=-\omega_{1}(\overline{a}_{1},\overline{a}_{2})=\int_{\Sigma}\bigl(\overline{a}_{1}a_{2}+a_{1}\overline{a}_{2}\bigr)=\omega_{2}(a_{1},a_{2}).$
@@ 2.8 Estimate for the family of the associated Laplacian operators
Let $X$ be a compact Riemann surface with a conformal metric $g_{X}$. Let $(E,\theta)$ be a generically regular semisimple Higgs bundle of degree 0 whose spectral curve is smooth. We fix a flat metric $h_{\det(E)}$ of $\det(E)$. For each $t>0$, there exists a harmonic metric $h_{t}$ of $(E,t\theta)$. We recall that the sequence $h_{t}$ is convergent to the limiting configuration $h_{\infty}$ in the $C^{\infty}$-sense locally on $X^{\circ}$.
Let $\mathop{\rm End}\nolimits_{0}(E)$ denote the trace free part of $\mathop{\rm End}\nolimits(E)$. Let $(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{t}}$ denote the formal adjoint of $\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta$ with respect to $h_{t}$ and $g_{X}$. We consider the operators $\Delta_{t}=(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{t}}\circ(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)$ on $A^{0}(\mathop{\rm End}\nolimits_{0}(E))$. For any sections $s$ of $A^{0}(\mathop{\rm End}\nolimits_{0}(E))$, let $\|s\|_{L^{2},h_{t}}$ denote the $L^{2}$-norm of $s$ with respect to $h_{t}$ and $g_{X}$.
@@
Proof  If $s\in A^{0}(\mathop{\rm End}\nolimits_{0}(E))$ satisfies $\Delta_{t}s=0$, it is standard to obtain $(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)s=0$. Because $(E,t\theta)$ is stable, we obtain $s=0$.
@@
By taking a subsequence, we may assume that the sequence $s_{i}$ is weakly convergent in $L_{1}^{2}$ locally on $X^{\circ}$. Let $s_{\infty}$ denote the limit. The sequence $s_{i}$ $(i=1,2,\ldots)$ is convergent to $s_{\infty}$ in $L^{p}$ for any $p>1$ locally on $X^{\circ}$. We obtain $\theta(s_{\infty})=0$ and $\overline{\partial}_{E}(s_{\infty})=0$. We also have $\|s_{\infty}\|_{L^{2},h_{\infty}}\leq 1$. Because $\mathop{\rm tr}\nolimits s_{i}=0$, we obtain $\mathop{\rm tr}\nolimits s_{\infty}=0$.
@@
Let $U$ be a relatively compact neighbourhood of $P_{\infty}$ in $X$ with a holomorphic coordinate $z$. Let $(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{i}}$ denote the formal adjoint of $\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta$ with respect to $h_{i}$ and the Euclidean metric $dz\,d\overline{z}$. By Lemma 2.44 below, we obtain
$\partial_{z}\partial_{\overline{z}}|\widetilde{s}_{i}|_{h_{i}}^{2}\leq 2\bigl|(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{i}}(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)(\widetilde{s}_{i})\bigr|_{h_{i}}\cdot|\widetilde{s}_{i}|_{h_{i}}=2a_{t(i)}g_{U}|\widetilde{s}_{i}|_{h_{i}}^{2}\leq 2a_{t(i)}g_{U}.$
@@
$\partial_{z}\partial_{\overline{z}}\bigl(|\widetilde{s}_{i}|_{h_{i}}^{2}-2a_{t(i)}\beta_{U}\bigr)\leq 0.$
@@
Let us consider the case $P_{\infty}\in D(\theta)$. Let $U^{\prime}$ be a relatively compact neighbourhood of $P_{\infty}$ in $U$ such that $U^{\prime}\cap D(\theta)=\{P_{\infty}\}$. By applying the maximum principle to the subharmonic function $|\widetilde{s}_{i}|_{h_{i}}^{2}-2a_{t(i)}\beta_{U}$, there exist $Q_{i}\in\partial U^{\prime}$ such that
@@
$\partial_{z}\partial_{\overline{z}}|s|_{h}^{2}=2\mathop{\rm Re}\nolimits\bigl((-\nabla_{z}\nabla_{\overline{z}}+f_{h}^{\dagger}\circ f)s,s\bigr)-|\nabla_{z}s|_{h}^{2}-|\nabla_{\overline{z}}s|_{h}^{2}-|f(s)|_{h}^{2}-|f^{\dagger}_{h}(s)|_{h}^{2}.$ (32)
@@
$\partial_{z}\partial_{\overline{z}}|s|_{h}^{2}=-h\bigl(\nabla_{z}\nabla_{\overline{z}}s,s\bigr)-h\bigl(s,\nabla_{\overline{z}}\nabla_{z}s\bigr)-\bigl|\nabla_{\overline{z}}s\bigr|_{h}^{2}-\bigl|\nabla_{z}s\bigr|_{h}^{2}\\ =-h\bigl(\nabla_{z}\nabla_{\overline{z}}s,s\bigr)-h\bigl(s,\nabla_{z}\nabla_{\overline{z}}s\bigr)-h\bigl(s,[f,f^{\dagger}_{h}]s\bigr)-\bigl|\nabla_{\overline{z}}s\bigr|_{h}^{2}-\bigl|\nabla_{z}s\bigr|_{h}^{2}\\ =2\mathop{\rm Re}\nolimits h\bigl(-\nabla_{z}\nabla_{\overline{z}}s,s\bigr)+2h\bigl(s,f^{\dagger}_{h}\circ f(s)\bigr)-\bigl|f(s)\bigr|_{h}^{2}-\bigl|f^{\dagger}_{h}(s)\bigr|_{h}^{2}-\bigl|\nabla_{\overline{z}}s\bigr|_{h}^{2}-\bigl|\nabla_{z}s\bigr|_{h}^{2}\\ =2\mathop{\rm Re}\nolimits h\bigl((-\nabla_{z}\nabla_{\overline{z}}+f^{\dagger}_{h}f)s,s\bigr)-\bigl|f(s)\bigr|_{h}^{2}-\bigl|f^{\dagger}_{h}(s)\bigr|_{h}^{2}-\bigl|\nabla_{\overline{z}}s\bigr|_{h}^{2}-\bigl|\nabla_{z}s\bigr|_{h}^{2}.$ (33)